OpenStudy (anonymous):

Which of the following is the largest volume? 1.2 x 101 m3 1.2 x 108 cm3 2.0 x 104 dm3 1.2 x 108 mm3

OpenStudy (somy):

to compare, convert all of them to 1 same unit

OpenStudy (somy):

lets take them all in mm3

OpenStudy (somy):

m3 * 10^9 = mm3

OpenStudy (somy):

cm3 * 1000= mm3

OpenStudy (somy):

dm3 *10^6 = mm3

OpenStudy (somy):

after converting them all to same unit u can see which one is the biggest number